OpenStudy (anonymous):

What is the sum of the geometric sequence 3, 15, 75, … if there are 8 terms?

OpenStudy (amistre64):

\[a\frac{1-r^n}{1-r}\] i believe

OpenStudy (amistre64):

r = 5 since its a geometric sequence and 15/3 = 5

OpenStudy (amistre64):

n=8 since it says there are 8 terms

OpenStudy (amistre64):

and a = 3 since that is the first term ....

OpenStudy (amistre64):

A = {3,15,75,...} 3 A = {1,5,25, ...} 3\(\sum\)A = \(3\cfrac{1-r^n}{1-r}\)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=3%281-5^8%29%2F%281-5%29 with any luck :)